OpenStudy (anonymous):

factor the expression completely 9x^2-1

OpenStudy (anonymous):

9x^2=1 x^2=1/9 "using the sq.root" x=1/3

OpenStudy (anonymous):

thats not and equals sign its subtraction im looking for (ax+/- b)(ax+/-b)

OpenStudy (anonymous):

oh srry i misunderstood